OpenStudy (anonymous):

first answer gets a medal!!!!! 5. What is the area of the shaded part of the parallelogram? A. 32 m² B. 128 m² C. 184 m² D. 256 m²

OpenStudy (anonymous):

@mathmath333 anybody please!!!!!

OpenStudy (mathmath333):

area of the shaded portion will be = area of parallelogram - area of the square area of parallelogram\(\large\tt \color{black}{=h\times b}\) \(\large\tt \color{black}{=12\times 16}\) \(\large\tt \color{black}{=192}\) area of square\(\large\tt \color{black}{=a^2}\) \(\large\tt \color{black}{=8^2}\) \(\large\tt \color{black}{=64}\) area of the shaded portion will be = 192 - 64 \(\huge\tt \color{black}{=128}\)

OpenStudy (anonymous):

B

OpenStudy (anonymous):

@mathmath333 I gave you a medal